OpenStudy (anonymous):

f(x)=10x-12 and g(x)=x^2-x. evalute the following expression f(0)+g(4)=

OpenStudy (anonymous):

f (0) = 10*0-12 = -2 g(4) = 4^2-4 = 12 f(0)+g(4) = 10